# RTop-K: Ultra-Fast Row-Wise Top-K Algorithm and GPU Implementation for Neural Networks

### v1

@@ Abstract
Top-k algorithms are essential in various applications, from high-performance computing and information retrieval to big data and neural network model training. This paper introduces RTop-K, a highly efficient parallel row-wise top-k selection algorithm designed for GPUs. RTop-K employs a Binary Search-based approach to optimize resource allocation and provides a scalable solution that significantly accelerates top-k operations. We perform a theoretical analysis of the effects of early stopping in our algorithm, demonstrating that it maintains the accuracy of neural network models while enhancing performance. Comprehensive tests show that our GPU implementation of RTop-K outperforms other row-wise top-k GPU implementations, with minimal impact on testing accuracy when early stopping is applied. Notably, RTop-K achieves speed increases ranging from 4.245× to 9.506× with early stopping, and 3.936× without early stopping, compared to state-of-the-art implementations. The proposed methods offer significant improvements in the training and inference of Graph Neural Networks (GNNs), addressing critical challenges in latency and throughput on GPU platforms.
The implementation can be found on Github [^1] https://anonymous.4open.science/r/RTopK-4DE7.
Introduction
@@
Top-k selection is a classic algorithmic challenge that involves identifying the k largest or smallest elements from n input elements based on some predefined ranking criteria. The top-k selection algorithm has been widely applied in many traditional scenarios, such as high-performance computing (HPC) (Muneer 2021), information retrieval (IR) (Ding and Suel 2011), big data (Gaihre, Pandey, and Liu 2019), and data mining (Malkov and Yashunin 2018). Today, the top-k algorithm is increasingly applied in the training and inference of neural network models. For example, the Avg-TopK (Özdemir 2023) pooling method has achieved more successful results in image classification accuracy and transfer learning models compared to traditional methods. TopK-SGD (Shi et al. 2019) applied to gradient sparsification techniques significantly reduces the communication traffic without obvious impact on the model accuracy. Combining top-k with sparse training (Jayakumar et al. 2021) can maintain constant sparsity and perform well while reducing resource requirements. In a study (Cui et al. 2021), a top-k attention loss function was introduced to address the top-k ranking prediction problem.
@@
Graph Neural Networks (GNNs) have drawn tremendous attention in the past years due to their unique ability to extract latent information from graph data (Hu et al. 2020). Practical applications of GNNs include the prediction of cascading power-grid failures (Liu et al. 2020), traffic forecasting (Jiang and Luo 2022), recommendation systems (Wu et al. 2020), and drug discovery (Bongini et al. 2021). In the design and acceleration of GNN training and inference, GPU platforms have become the prevalent choice due to their multiple advantages. Firstly, compared to other processing hardware, GPUs provide superior processing power and memory throughput (Li et al. 2018). For example, the NVIDIA A6000 GPU boasts an impressive computation capability of 38.7 Tera FLOPS and a memory throughput 768 GB/s. Secondly, many leading supercomputers (such as Aurora (The Verge 2021) and El Capitan (Moss 2021)) use GPUs as their primary computing resource. Thirdly, many applications and services related to deep learning and neural networks are developed and deployed on GPU platforms. However, GNN training and inference still typically pose strict challenges on latency and throughput (Xie et al. 2023).
@@
Figure 1: The core operation of MaxK-GNN, which introduces top-k selection into the GNN workflow to provide non-linearity and acceleration.
@@
Recently, MaxK-GNN (Peng et al. 2024) has achieved great success in the acceleration and optimization of GNN training and inference on GPUs. As shown in Figure 1, this work introduces top-k selection before the feature aggregation step in GNNs, which not only provides nonlinearity in GNNs to optimize the model’s expressive ability but also demonstrates that performing SPMM operations in GNNs with the top-k-processed right multiplication matrix can achieve several times speedup over traditional workflows while maintaining good model accuracy. The top-k selection operation in Max-GNN necessitates performing a large-scale row-wise top-k computation, i.e., executing top-k operations simultaneously across a batch of vectors on GPUs.
@@
Traditional top-k algorithms and their GPU implementations (Gaihre et al. 2021; Zhang et al. 2023; Li et al. 2024) are typically optimized for single queries or limited batched queries, that is, for a large vector or a small batch of large vectors (typically with a batch size not exceeding 100). However, the optimization focus for traditional scenarios differs from the row-wise top-k algorithm required for GNN training and inference. Implementing and optimizing row-wise top-k algorithms on GPUs pose challenges in terms of dispersion, parallelism, and efficiency. Since row-wise top-k involves performing top-k operations on a large batch of vectors simultaneously, and each vector’s length corresponds to the hidden dimension of the neural network layer (which usually does not exceed 1024), it is crucial to allocate only a small and appropriate amount of GPU resources for each vector. Under these limited resource constraints, the various optimization methods proposed for large vectors in traditional top-k implementations may be overly complex and inefficient. We should seek simple and efficient algorithms tailored to this scenario.
@@
  • We propose a Binary Search-based Top-k Algorithm and provide a theoretical analysis of the effects of early stopping.
  • We implement the Binary Search-based Top-k Algorithm on GPU and conduct comprehensive tests, demonstrating that it outperforms state-of-the-art row-wise top-k GPU implementations, with early stopping having minimal impact on the model’s testing accuracy.
Preliminary and Related Works
Top-k Algorithms
The heap-based top-k algorithm (Cormen et al. 2009) uses a min-heap to maintain the top-k elements. By iterating through the data, each element is compared with the smallest element in the heap (i.e., the root of the heap). If the current element is larger, it replaces the root, and the heap is restructured to maintain the min-heap property. On the other hand, if the k smallest elements are required, a max-heap is used instead. QuickSelect (Dashti, Komarov, and D’Souza 2013) is a top-k algorithm inspired by the quicksort algorithm, designed to find the k-th largest element in an unordered list. It works by partitioning the data around a pivot element and recursively processing the part that contains the k-th element. The bucket-based top-k algorithm (Alabi et al. 2012) divides the data into several buckets based on the range of values. The top-k elements are then found by sorting the buckets or partially sorting only the relevant buckets. This method is particularly useful for uniformly distributed data. RadixSelect (Alabi et al. 2012) is a variant of radix sort used for selecting the top-k elements. It processes the digits of the numbers starting from the least significant digit to the most significant digit. This method is efficient for fixed-length integer keys. The bitonic top-k algorithm (Shanbhag, Pirk, and Madden 2018) is based on bitonic sorting, a parallel sorting algorithm. It constructs a bitonic sequence (a sequence that first increases and then decreases) and then merges it to find the top-k elements.
@@
Figure 2: Illustration of the Binary Search-based Top-k Algorithm
GPU Architecture
@@
GPU Top-k Implementations
Dr. Top-k (Gaihre et al. 2021) is a delegate-centric system that helps reduce the workload of GPU top-k computations, including Radix Select, Bucket Select, and Bitonic Select. It achieves this by dividing the input into sub-ranges and selecting delegates from them, along with performing multi-GPU optimizations. A work (Zhang et al. 2023) proposed two optimization methods, AIR Top-k and GridSelect. AIR Top-k employs an iteration-fused design and adaptive strategy to minimize CPU-GPU communication and memory traffic, while GridSelect uses a shared queue and parallel two-step insertion to decrease costly operations, enhancing parallel top-k efficiency on GPUs. A recent RadixSelect implementation RadiK (Li et al. 2024) utilizes an optimization framework tailored for high memory bandwidth and resource utilization, along with an adaptive scaling technique for enhanced robustness, that supports larger k values with high efficiency.
@@
PyTorch’s top-k implementation (Pytorch 2024) is suitable for row-wise top-k operations. However, it involves a precise sorting operation for each vector, which can lead to lower efficiency. Despite being able to handle large batch sizes, the exact sorting required for each vector results in a performance bottleneck, making it less efficient for large-scale applications where rapid top-k selection is critical.
RTop-K Framework
Algorithm 1 Binary Search-based Top-k Algorithm
@@
4:  $mid,cnt\leftarrow 0$
5:  while $max-min>\epsilon$ do
6:     $mid\leftarrow\frac{min+max}{2}$
7:     $cnt\leftarrow\left|\{i\mid\mathbf{v}_{i}\geq mid\}\right|$
8:     if $cnt<k$ then
9:        $max\leftarrow mid$
10:     else if $cnt>k$ then
11:        $min\leftarrow mid$
12:     else
13:        break
14:     end if
@@
16:  if $cnt=k$ then
17:     $\mathbf{elems},\mathbf{indices}\leftarrow\{(\mathbf{v}_{i},i)\mid\mathbf{v}_{i}\geq mid\}$
18:  else
19:     $\mathbf{elems},\mathbf{indices}\leftarrow\{(\mathbf{v}_{i},i)\mid\mathbf{v}_{i}>mid+\epsilon\}$
20:     Append the first $k-\left|\mathbf{elems}\right|$ pairs of $\{(\mathbf{v}_{i},i)\mid mid-\epsilon\leq\mathbf{v}_{i}\leq mid+\epsilon\}$ to $\mathbf{elems},\mathbf{indices}$
21:  end if
22:  return  $\mathbf{elems},\mathbf{indices}$
@@
Avg 7.60 8.29 8.95 9.52 9.60
Algorithm 2 Binary Search-based Top-k Algorithm with Early Stopping
@@
4:     $mid\leftarrow\frac{min+max}{2}$
5:     $cnt\leftarrow\left|\{i\mid\mathbf{v}_{i}\geq mid\}\right|$
@@
7:        $max\leftarrow mid$
8:     else if $cnt>k$ then
9:        $min\leftarrow mid$
10:     else
11:        break
12:     end if
13:  end for
14:  $\mathbf{elems},\mathbf{indices}\leftarrow\{(\mathbf{v}_{i},i)\mid\mathbf{v}_{i}\geq min\}$
15:  return  first $k$ pairs of $\mathbf{elems},\mathbf{indices}$
Binary Search-based Top-k Algorithm
The algorithm, as illustrated in Fig. 2, first retrieves the $min$ and $max$ values of the vector, and then uses several iterations of binary search to determine a threshold. The algorithm stops when the number of elements filtered by the current threshold equals $k$. To address the issue where the loop might struggle to exit due to multiple equal or very close elements during filtering, we introduce the condition $max-min>\epsilon$, where $\epsilon=\epsilon^{\prime}\cdot max$, and $\epsilon^{\prime}$ is a small value representing the precision, such as $10^{-4}$. If this condition is not met, the loop will also exit, and the remaining elements will be selected sequentially among those that are equal within the specified precision. Table 1 presents the statistical results of the iteration counts at which the algorithm exits for different values of $k$, with the vector’s size $M=256$. For each $k$ value, $10^{5}$ repeated experiments were conducted, with the vector initialized with normally distributed elements. It can be observed that the average iteration at exit ranges from $7.6$ to $9.6$, and the probability of the iteration count being less than or equal to 13 exceeds $95\%$. Algorithm 1 summarizes the complete binary search-based top-k algorithm process, but it still contains a number of branching conditions. Given the inherent robustness of neural networks, we can explore the feasibility of incorporating early stopping into the algorithm. We first present the pseudocode for the early stopping algorithm and then conduct a numerical analysis.
As shown in Algorithm 2, the introduction of early stopping further simplifies the algorithm, with the main loop forcefully exiting in no more than $max\_iter$ iterations. The collection phase uses $min$ instead of $mid$ as the threshold, ensuring that only one-pass collection is needed, thereby eliminating the need for the two-pass collection process present in the original algorithm (in the branch where $cnt\neq k$). Table 2 summarizes the hit rate and differences between the early stopping top-k selection with different $max\_iter$ values and the optimal top-k selection. The experiments were conducted with vectors of size $M=256$ consisting of normally distributed elements, and $10^{5}$ repeated experiments for each condition. When $max\_iter\geq 5$ for $k\geq 32$ ($max\_iter\geq 6$ for $k=16$), both the maximum element and the minimum element in the early stopping top-k selection have an average relative error of no more than 5%. For $k\geq 64$, only 4 iterations are needed for the hit rate between the early stopping top-k selection and the optimal top-k selection to exceed 80%. These results indicate that the early stopping top-k selection is numerically stable and controllable. We will further test the impact of early stopping top-k selection on model accuracy in the experimental section.
Figure 3: Example of MaxK hardware implementation
@@
GPU Implementation
@@
Searching stage: In this stage, each vector is handled by a single warp, assuming the warp contains $w$ threads (Fig. 3 illustrates $w=4$, while in actual hardware environments $w=32$). The first step is to obtain the maximum and minimum elements of the vector. The vector is first divided into $\lceil M/w\rceil$ parts, with each thread responsible for extracting the maximum and minimum elements within its assigned part. Then, a tree-reduction using the shuffle primitive is performed in five steps to obtain the maximum and minimum elements across the entire warp, and these values are broadcasted to all threads within the warp. The second step is to perform binary search according to Algorithm 2 to obtain the selection threshold. In each iteration, the count of elements above the current threshold is accumulated and broadcasted using the same approach. After a specified number of iterations, the final threshold is obtained.
@@
Experiments
Figure 4: Testing accuracy and overall speed-up ratio of applying RTop-K to GraphSage training on different graphs.
@@
Figure 5: Comparison of execution time (ms) between RTop-K with different early stopping $max\_iter$ and without early stopping ($\epsilon=10^{-4}$), against PyTorch for various ($N,M,k$) configurations.
Setup and Configuration
The CUDA source code of RTop-K of the Binary Search-based Top-k Algorithm with Early Stopping is compiled utilizing NVCC, version 12.2, and the execution is carried out on an NVIDIA A6000 platform equipped with Ubuntu 22.04. The tests cover various input matrix dimensions, with the number of rows $N$ ranging from $2^{14}$ to $2^{20}$, hidden dimensions $M$ ranging from 256 to 768, and $k$ values ranging from 16 to 128.

### v3

# RTop-K: Ultra-Fast Row-Wise Top-K Selection for Neural Network Acceleration on GPUs

@@ Abstract
Top-k selection algorithms are fundamental in a wide range of applications, including high-performance computing, information retrieval, big data processing, and neural network model training. In this paper, we present RTop-K, a highly efficient parallel row-wise top-k selection algorithm specifically designed for GPUs. RTop-K leverages a binary search-based approach to optimize row-wise top-k selection, providing a scalable and accelerated solution. We conduct a detailed analysis of early stopping in our algorithm, showing that it effectively maintains the testing accuracy of neural network models while substantially improving performance. Our GPU implementation of RTop-K demonstrates superior performance over state-of-the-art row-wise top-k GPU implementations, achieving an average speed-up of up to 11.49×11.49$\times$11.49 × with early stopping and 7.29×7.29$\times$7.29 × without early stopping. Moreover, RTop-K accelerates the overall training workflow of MaxK-GNNs, delivering speed-ups ranging from 11.97% to 33.29% across different models and datasets.
The GPU implementation can be found on Github [^2] https://github.com/xiexi51/RTopK.
11footnotetext: These authors contributed equally. +Corresponding author.
1 Introduction
@@
Top-k selection is a classic algorithmic challenge that involves identifying the k largest or smallest elements from n input elements based on some predefined ranking criteria. The top-k selection algorithm has been widely applied in many traditional scenarios, such as high-performance computing (HPC) (Muneer, 2021), information retrieval (IR) (Ding & Suel, 2011), big data (Gaihre et al., 2019), and data mining (Malkov & Yashunin, 2018). Today, the top-k algorithm is increasingly applied in the training and inference of neural network models. For example, the Avg-TopK (Özdemir, 2023) pooling method has achieved more successful results in image classification accuracy and transfer learning models compared to traditional methods. TopK-SGD (Shi et al., 2019) applied to gradient sparsification techniques significantly reduces the communication traffic without obvious impact on the model accuracy. Combining top-k with sparse training (Jayakumar et al., 2021) can maintain constant sparsity and perform well while reducing resource requirements. In a study (Cui et al., 2021), a top-k attention loss function was introduced to address the top-k ranking prediction problem.
@@
Graph Neural Networks (GNNs) have drawn tremendous attention in the past years due to their unique ability to extract latent information from graph data (Hu et al., 2020). In the design and acceleration of GNN training and inference, GPU platforms have become the prevalent choice due to their multiple advantages. Firstly, compared to other processing hardware, GPUs provide superior processing power and memory throughput (Li et al., 2018). For example, the NVIDIA A6000 GPU boasts an impressive computation capability of 38.7 Tera FLOPS and a memory throughput 768 GB/s. Secondly, many leading supercomputers (such as Aurora and Eagle (LiveScience, 2023)) use GPUs as their primary computing resource. Thirdly, many applications and services related to deep learning and neural networks are developed and deployed on GPU platforms. However, GNN training and inference still typically pose strict challenges on latency and throughput (Xie et al., 2023).
@@
Figure 1: The core operation of MaxK-GNN, which introduces row-wise top-k selection into the GNN workflow to provide non-linearity and acceleration.
@@
Recently, MaxK-GNN (Peng et al., 2024) has achieved great success in the acceleration and optimization of GNN training and inference on GPUs. As shown in Figure 1, this work introduces row-wise top-k selection before the feature aggregation step in GNNs, which not only provides nonlinearity in GNNs to optimize the model’s expressive ability but also demonstrates that performing SPMM operations in GNNs with the row-wise top-k-processed right hand matrix can achieve several times speedup over traditional workflows while maintaining good model accuracy. The top-k selection operation in Max-GNN necessitates performing a large-scale row-wise top-k computation, i.e., executing top-k operations simultaneously across a batch of vectors on GPUs.
@@
Traditional top-k algorithms and their GPU implementations (Gaihre et al., 2021; Zhang et al., 2023; Li et al., 2024) are typically optimized for single queries or limited batched queries, that is, for a large vector or a small batch of large vectors (typically with a batch size not exceeding 100). However, the optimization focus for traditional scenarios differs from the row-wise top-k algorithm required for GNN training and inference. Implementing and optimizing row-wise top-k algorithms on GPUs pose challenges in terms of dispersion, parallelism, and efficiency. Since row-wise top-k involves performing top-k operations on a large batch of vectors simultaneously, and each vector’s length corresponds to the hidden dimension of the neural network layer (which usually does not exceed 1024), it is crucial to allocate only a small and appropriate amount of GPU resources for each vector. Under these limited resource constraints, the various optimization methods proposed for large vectors in traditional top-k implementations may be overly complex and inefficient. We should seek simple and efficient algorithms tailored to this scenario.
@@
  • We provide a comprehensive summary of GPU top-k selection algorithms and analyze the performance limitations of state-of-the-art GPU implementations in the row-wise top-k selection scenario.
  • We propose a binary search-based top-k selection algorithm and provide a theoretical analysis of the effects of early stopping.
  • We implement the binary search-based top-k selection algorithm on the GPU and conduct comprehensive tests, demonstrating that it outperforms state-of-the-art row-wise top-k GPU implementations, with early stopping having minimal impact on testing accuracy.
2 Preliminary and Related Works
2.1 Top-k Algorithms
The heap-based top-k algorithm (Cormen et al., 2009) uses a min-heap to maintain the top-k elements by comparing and replacing the heap root when a larger element is encountered. QuickSelect (Dashti et al., 2013) leverages a partition-based approach similar to quicksort to find the k-th largest element. The bucket-based algorithm (Yang et al., 2024) divides data into buckets, sorting only relevant ones to find the top-k elements, which is effective for uniformly distributed data. RadixSelect (Alabi et al., 2012) uses digit-wise sorting to identify the top-k elements efficiently. The bitonic top-k algorithm (Shanbhag et al., 2018) employs bitonic sorting to merge and find the top-k elements in parallel.
@@ 2.1 Top-k Algorithms
2.2 GPU Architecture
@@
2.3 GPU Top-k Implementations
Dr. Top-k(Gaihre et al., 2021) is a delegate-centric system that helps reduce the workload of GPU top-k computations, including Radix Select, Bucket Select, and Bitonic Select. It achieves this by dividing the input into sub-ranges and selecting delegates from them, along with performing multi-GPU optimizations. A work(Zhang et al., 2023) proposed two optimization methods, AIR Top-k and GridSelect. AIR Top-k employs an iteration-fused design and adaptive strategy to minimize CPU-GPU communication and memory traffic, while GridSelect uses a shared queue and parallel two-step insertion to decrease costly operations, enhancing parallel top-k efficiency on GPUs. A recent RadixSelect implementation RadiK(Li et al., 2024) utilizes an optimization framework tailored for high memory bandwidth and resource utilization, along with an adaptive scaling technique for enhanced robustness, that supports larger k values with high efficiency.
@@ 2.3 GPU Top-k Implementations
PyTorch’s top-k implementation (Pytorch, 2024) is suitable for row-wise top-k operations. It uses RadixSelect as the underlying method, which, as analyzed in Section 2.1, is overly complex for each limited-size vector. Moreover, its selection results are sorted, which is also unnecessary. Although it can handle large batch sizes, its efficiency is suboptimal in scenarios where a minimalistic top-k selection is critical for each single vector.
Algorithm 1 Binary Search-based Top-k Selection Algorithm
@@
4:  $thres,cnt\leftarrow max,0$
5:  while $thres-min>\epsilon$ do
6:     $cnt\leftarrow\left|\{i\mid\mathbf{v}_{i}\geq thres\}\right|$
7:     if $cnt<k$ then
8:        $max\leftarrow thres$
9:     else if $cnt>k$ then
10:        $min\leftarrow thres$
11:     else
12:        break
13:     end if
14:     $thres\leftarrow\frac{min+max}{2}$
@@
16:  $borderline\_cnt\leftarrow k-cnt$
17:  if $borderline\_cnt\geq 0$ then
18:     $\mathbf{elems},\mathbf{indices}\leftarrow\{(\mathbf{v}_{i},i)\mid\mathbf{v}_{i}\geq thres\}$
19:     Append the first $borderline\_cnt$ pairs of $\{(\mathbf{v}_{i},i)\mid min\leq\mathbf{v}_{i}<thres\}$ to $\mathbf{elems},\mathbf{indices}$
20:  else
21:     $\mathbf{elems},\mathbf{indices}\leftarrow$ the first $k$ pairs of $\{(\mathbf{v}_{i},i)\mid\mathbf{v}_{i}\geq thres\}$
22:  end if
23:  return  $\mathbf{elems},\mathbf{indices}$
3 RTop-K Framework
3.1 Binary Search-based Top-k Selection Algorithm
The algorithm, as illustrated in Fig. 2, first retrieves the $min$ and $max$ values of the vector, and then uses several iterations of binary search to determine a threshold $thres$. The $min$, $max$, and $thres$ values are updated in each iteration, and the loop terminates when the number of elements filtered by the current threshold equals $k$. One corner case that must be considered is when $\mathbf{v}$ contains multiple equal or very close elements, and these elements happen to be near the ”borderline” during top-k selection. In this scenario, relying solely on $thres$ for selection makes it difficult or even impossible to extract exactly $k$ elements. It can be verified that in such cases, these borderline elements are always located between $min$ and $thres$. Therefore, a second filtering step using $min$ is sufficient to supplement the selection and ensure that exactly $k$ elements are chosen. The parameter $\epsilon$ represents the precision of the algorithm, which determines the maximum width of the borderline and imposes an upper bound on the number of iterations in the loop. When $\epsilon$ is set to 0, Algorithm 1 becomes an exact top-k selection algorithm.
Figure 2: Illustration of the binary search-based top-k selection algorithm.
@@ 3.1 Binary Search-based Top-k Selection Algorithm
Average Exit 7.60 8.29 8.95 9.52 9.60
Algorithm 2 Binary Search-based Top-k Selection Algorithm with Early Stopping
@@
4:     $thres\leftarrow\frac{min+max}{2}$
5:     $cnt\leftarrow\left|\{i\mid\mathbf{v}_{i}\geq thres\}\right|$
@@
7:        $max\leftarrow thres$
8:     else
9:        $min\leftarrow thres$
10:     end if
11:  end for
12:  $\mathbf{elems},\mathbf{indices}\leftarrow\{(\mathbf{v}_{i},i)\mid\mathbf{v}_{i}\geq min\}$
13:  return  first $k$ pairs of $\mathbf{elems},\mathbf{indices}$
Table 1 presents the statistical results of the iteration counts at which the algorithm exits for different values of $k$, with the vector’s size $M=256$. For each $k$ value, $10^{5}$ repeated experiments were conducted, with the vector initialized with normally distributed elements. It can be observed that the average iteration at exit ranges from $7.6$ to $9.6$, and the probability of the iteration count being less than or equal to 13 exceeds $95\%$.
Algorithm 1 summarizes the complete binary search-based top-k algorithm process. It contains a number of branching conditions, and the loop length executed by each warp can be different. We attempt to further simplify it.
Given the inherent robustness of neural networks, we can explore the feasibility of incorporating early stopping into the algorithm. We present the pseudocode for the early stopping algorithm and then conduct a numerical analysis.
As shown in Algorithm 2, the introduction of early stopping further simplifies the algorithm, with the main loop forcefully exiting in no more than $max\_iter$ iterations. The collection phase uses $min$ as the threshold, ensuring that only one-pass collection is needed. Table 2 summarizes the hit rate (overlap ratio) and the average relative error between the early stopping top-k selection with different $max\_iter$ settings and the optimal top-k selection. The experiments were conducted with vectors of size $M=256$ consisting of normally distributed elements, and $10^{5}$ repeated experiments for each condition. When $max\_iter\geq 5$ for $k\geq 32$ ($max\_iter\geq 6$ for $k=16$), both the maximum element and the minimum element in the early stopping top-k selection have an average relative error of no more than 5%. For $k\geq 64$, only 4 iterations are needed for the hit rate between the early stopping top-k selection and the optimal top-k selection to exceed 80%. These results indicate that the early stopping top-k selection is numerically stable and controllable. We will further test the impact of early stopping top-k selection on model accuracy in the experimental section.
@@
Figure 3: GPU implementation of the binary search-based top-k selection algorithm.
3.2 GPU Implementation Design
@@
Searching stage: In this stage, each vector is handled by a single warp, assuming the warp contains $w$ threads (Fig. 3 illustrates $w=4$, while in actual hardware environments $w=32$). The key point of the implementation is that the $max$, $min$, and $thres$ values at the beginning and in each iteration need to be synchronized across all threads within the warp. This can be achieved using a combination of classic tree-reduction and broadcast primitives. The first step is to obtain the $max$ and $min$ of the vector. The vector is divided into $\lceil M/w\rceil$ parts, with each thread responsible for extracting the maximum and minimum elements within its assigned part. Then, a tree-reduction using the shuffle primitive is performed in $\log_{2}{w}$ steps to obtain the maximum and minimum elements across the entire warp, and these values are broadcasted to all threads within the warp. The second step is to perform a binary search according to Algorithm 2 to obtain the selection threshold. In each iteration, the count of elements above the current threshold is accumulated and broadcasted using the same tree-reduction method. Then, each thread uses $cnt$ to simultaneously update the $max$, $min$, and $thres$ values. After a specified number of iterations, the final threshold is obtained.
@@
Figure 4: Comparison of kernel execution time (ms) between RTop-K with different early stopping $max\_iter$ values and without early stopping ($\epsilon=10^{-16}$), against PyTorch for various configurations of ($N$, $M$, $k$), where $N=2^{14},2^{16},2^{18},2^{20}$, $M=256,512,768$, and $k=16,32,64,96,128$. The average speedup of the no early stopping version for each $(N,M)$ setting is indicated in the title of each subplot.
4 Experiments
4.1 Setup and Configuration